# Enhancing Biomedical Knowledge Discovery for Diseases: An Open-Source Framework Applied on Rett Syndrome and Alzheimer’s Disease

### v2

@@ Appendix B Annotation Portal
Figure 7 presents the annotation portal with an example from the ReDReS dataset. The annotator’s task is to identify the semantic relation between the two highlighted entities, classifying it as either a Positive Relation, Negative Relation, Complex Relation, or No Relation. If the sentence is considered uninformative or if there are errors in entity detection, type, or span, the annotator can remove the sentence or the entities. Furthermore, the annotator is encouraged to provide feedback, including any additional text that can clarify or elaborate on the relationship between the entities. By providing this supplementary information, annotators can contribute to a richer and more nuanced understanding of the relations within the data.
Figure 7: Annotation portal: The annotator should define the semantic relation (Positive Relation, Negative Relation, Complex Relation, and No Relation) between the two highlighted entities: Rett syndrome and MECP2 gene. The annotator can remove the sentence if it is not informative and the entities if the entity detection or the type/span is incorrect. Additionally, the annotator can provide feedback, adding the text that is useful to define the relation between the entities.
@@ Appendix C Dataset Instances
  • These techniques have thus provided morphological and functional brain alterations mapping of Alzheimer’s disease: on one hand grey matter atrophy first concerns the medial temporal lobe before extending to the temporal neocortex and then other neocortical areas; on the other hand, metabolic alterations are first located within the posterior cingulate cortex and then reach the temporo-parietal area as well as the prefrontal cortex, especially in its medial part.
@@ Appendix D Localized Context Vector
Data Type1 Binary setup Multi-class setup
Micro Evaluation Macro Evaluation
ReDReS A 90.72/89.95 91.15/89.9 90.15/88.83 74.49/73.91 74.06/75.13 72.22/72.65 74.52/74.5 74.11/75.09 71.8/72.95
E 89.57/89.39 89.98/90.28 89.49/87.19 73.73/75.67 73.22/ 70.17/70.98 73.95/74.9 73.38/ 69.97/70.69
F 90.48/89.09 90.72/89.84 89.68/88.35 72.86/74.18 74.22/ 71.36/73.57 72.62/72.82 74.13/ 72/73.82
G 90.78/89.32 90.57/89.57 89.93/88.41 74.33/73.35 74/ 72.68/73.14 74.57/73.78 74.32/ 72.48/73.43
O 90.9/89.94 90.75/90.12 89.51/87.82 73.99/74.51 74/ 67.25/68.67 73.83/74.71 73.69/ 67.43/67.75
ReDAD A 88.31/90.15 88.73/89.18 87.49/89.19 77.64/77.07 78.4/76.38 76.04/75.96 77.34/77.21 77.68/76.66 76.45/75.62
D 87.33/88.99 88.74/89.72 87.56/89.63 78.28/76.73 78.27/77.03 76.49/76.02 78.26/76.8 77.89/77.18 76.25/76.04
E 88.03/88.63 88.43/90.48 87.71/87.86 77.83/77.45 78.96/78.02 75.35/73.65 77.75/77.34 78/78.1 75.36/73.28
F 87.71/89.54 88.62/89.86 87.63/88.94 77.59/76.5 77.43/77.25 76.34/75.93 77.35/76.33 77.48/77.45 75.47/76.1
G 88.17/90.06 88.72/90.38 87.55/88.85 77.83/77.64 77.76/76.99 76.35/75.71 77.5/77.69 77.33/76.41 76.61/75.73
O 88.27/90.59 88.58/89.91 86.19/87.74 77.08/79.02 76.89/74.85 /71.32 76.78/78.95 76.62/74.93 70.57/70.75
@@
Table 5: LaMReDA Results (%) in binary and multi-class setup utilizing BiomedBERT-base, BioLinkBERT-base, and BioGPT as backbone language model: Each cell shows the average F1-score from 10 runs (original test set) and from 5-fold cross-validation setup.
Appendix E Experiments with BioLinkBERT and BioGPT
Assessing the impact of different language models on the performance of the LaMReDA and LaMEL models, we conduct experiments using additional models. Specifically, we evaluate the effect of BioLinkBERT (Yasunaga et al., 2022) (base model) and BioGPT (Luo et al., 2022b), both pre-trained on PubMed abstracts. BioLinkBERT is selected to determine whether leveraging the links between documents in the pre-training corpus benefits the relation extraction task of our study. This model uses citations between PubMed articles to model document connections, which may enhance its ability to capture semantic relations. BioGPT, as the GPT-based (Radford et al., 2018, 2019) transformer decoder (Vaswani et al., 2017) representative, is incorporated to explore the generative model’s capabilities in capturing semantic relations between entities. Yasunaga et al. (2022) state that the pre-training data of BioLinkBERT and BioMedBERT are identical  [^14] https://pubmed.ncbi.nlm.nih.gov. Papers published before Feb. 2020., except that they also use the citations between PubMed articles. Luo et al. (2022b) clarify that BioGPT’s pre-training includes PubMed abstracts updated before 2021. Hence, all language models in our study are pre-trained on identical or nearly identical datasets, making the comparison between them reliable.
We conduct experiments using a variety of representations, specifically testing the $R_{A}$, $R_{D}$, $R_{E}$, $R_{F}$, $R_{G}$, and $R_{O}$ (Eq. 9, 12, 13, 14, 15, 23) relation representations of LaMReDA, as well as the $E_{A}$, $E_{D}$, $E_{E}$, and $E_{H}$ (Eq. 1, 4, 5, 8) entity representations of LaMEL. The experimental setup follows the same procedure outlined in subsection 3.3, and the experiments are performed on both the original split (Tab.1) and in a 5-fold cross-validation setting.
The results in Tables 5 and 6 demonstrate the F1-scores of the LaMReDA and LaMEL models on the ReDReS and ReDAD datasets, using different language models. The experiments show that both LaMReDA and LaMEL perform consistently well across various relations ($R_{A}$, $R_{D}$, $R_{E}$, $R_{F}$, $R_{G}$, $R_{O}$) and entities ($E_{A}$, $E_{D}$, $E_{E}$, $E_{H}$) representations, utilizing the different language models. This indicates the models’ ability to learn meaningful representations for the semantic relation detection task, independent of the initial token selection.
@@ Appendix E Experiments with BioLinkBERT and BioGPT
BiomedBERT2 BioLinkBERT2 BioGPT BiomedBERT2 BioLinkBERT2 BioGPT
A 90.25/89.43 90.68/89.02 89.43/ 86.73/88.75 87.71/ 86.14/
D 90.47/88.9 90.35/89.3 88.52/ 86.29/88.88 87.35/ 85.28/
E 90.61/89.1 90.25/89.42 89.48/ 86.03/88.96 87.05/ 87.42/
H 89.68/89.24 90.73/89.48 89.21/ 87.29/89.91 87.56/ 85.55/
@@
Table 6: LaMEL Results (%) in binary setup utilizing BiomedBERT-base, BioLinkBERT-base, and BioGPT as backbone language model: Each cell shows the average F1-score from 10 runs (original test set) and from 5-fold cross-validation setup.
An intra-lm comparison reveals that the use of relation representation $R_{O}$ results in lower performance across both datasets when BioGPT is employed, particularly in the multi-class setup (Tab. 5). This suggests that the context vector extracted from the generative model is less informative for this task. Conversely, when the context vector is derived from BiomedBERT and BioLinkBERT, the models show strong performance, showing the best results on ReDAD in the 5-fold cross-validation setup (Tab. 5).
The inter-language model comparison indicates that BiomedBERT and BioLinkBERT perform similarly, suggesting that leveraging document links during pre-training, as BioLinkBERT does, does not offer a significant advantage for our task. In contrast, the use of BioGPT leads to lower performance, particularly in the multi-class setup, emphasizing the superiority of encoder-based language models such as BiomedBERT and BioLinkBERT over decoder-based BioGPT for the relation detection task. Despite BioGPT’s larger size, with 347 million parameters (Luo et al., 2022b) compared with the 110 million parameters of BiomedBERT and BioLinkBERT (Vaswani et al., 2017; Yasunaga et al., 2022), the generative approach appears less effective for this particular semantic relation task, further underscoring the advantages of encoder-focused models in this context.
Appendix F Baseline Performance: Lower Bound
To establish a baseline performance (lower bound) for comparison, contrasting with the human evaluation that serves as an upper bound, we randomly assign class labels to each instance of the test set based on the prior class distribution in the training set (Tab. 1). This simulates a classifier with no ability to learn relations between entities. We repeat this experiment 1 million times for robustness and report the average F1-score.
Appendix G Error Analysis
Whereas the macro and micro evaluations in the multi-class setup indicate strong overall performance across different relation types, we conduct an error analysis to provide additional qualitative insights. Using the best-performing models on the ReDReS and ReDAD datasets based on the macro evaluation (Tab. 3), we extract predictions from the 5-fold cross-validation setup. Confusion matrices for each class are generated to present a comprehensive view of the misclassified instances.
Figure 8: ReDReS Confusion Matrices for each relation class using the best-performing model with BiomedBERT-base as the backbone language model.
Figure 9: ReDReS Confusion Matrices for each relation class using the best-performing model with BiomedBERT-large as the backbone language model.
Figure 10: ReDAD Confusion Matrices for each relation class using the best-performing model with BiomedBERT-base as the backbone language model.
Figure 11: ReDAD Confusion Matrices for each relation class using the best-performing model with BiomedBERT-large as the backbone language model.
We analyze the false negatives for each relation class to gain insights from the misclassification rates. In the ReDReS dataset, the majority of misclassified instances of the positive relation (324 and 331 instances using BiomedBERT-base and BiomedBERT-large, respectively) are labeled as complex relations (Tab. 7). This is intuitively reasonable, as the model recognizes that the entities are semantically related but fails to predict the correct relation type. Conversely, complex relation instances are often mislabeled as no relation (Tab. 8), highlighting the challenge of detecting nuanced semantic relations that require deeper reasoning. For the no relation class (Tab. 10), most false negatives (364 and 303 instances using BiomedBERT-base and BiomedBERT-large, respectively) are also labeled as complex relations, indicating a model bias toward predicting associations between co-occurring entities. Similar patterns are observed in the ReDAD dataset regarding the positive, complex, and no relation classes. The under-represented negative class has the most misclassified cases labeled as complex or positive relations in ReDReS and ReDAD, respectively, suggesting that while the model detects the semantic connection, it tends to predict one of the two more represented classes in the datasets.
Data BiomedBERT-base BiomedBERT-large
Complex Negative No Relation Complex Negative No Relation
ReDReS 331 14 41 324 11 42
ReDAD 424 27 40 425 21 18
Table 7: False Negative analysis for the positive relation type.
Data BiomedBERT-base BiomedBERT-large
Positive Negative No Relation Positive Negative No Relation
ReDReS 221 26 307 188 8 249
ReDAD 335 19 338 274 6 297
Table 8: False Negative analysis for the complex relation type.
Data BiomedBERT-base BiomedBERT-large
Positive Complex No Relation Positive Complex No Relation
ReDReS 25 28 1 15 34 1
ReDAD 44 23 2 44 25 1
Table 9: False Negative analysis for the negative relation type.
Data BiomedBERT-base BiomedBERT-large
Positive Complex Negative Positive Complex Negative
ReDReS 28 303 2 41 364 0
ReDAD 52 653 1 50 607 1
Table 10: False Negative analysis for the no relation relation type.
Appendix H Distantly Supervised Datasets
ReDReS and ReDAD include gold annotations for a small fraction of the extracted sentences. The pre-processed text consists of 28,622 and 1,301,429 additional sentences related to RS and AD respectively, without annotations about the semantic relation between the detected entities. Observing that the supervised models achieve performance levels comparable to human experts, we leverage the best-performing models to generate silver labels for the unannotated instances. For the binary setup, we employ:
  • LaMReDA (BiomedBERT large) with the relation representation $R_{A}$ (Eq. 9) for the RS corpus.
  • LaMReDA (BiomedBERT large) with the relation representation $R_{J}$ (Eq. 18) for the AD corpus.
For the multi-class setup, we use:
  • LaMReDA (BiomedBERT large) with the relation representation $R_{F}$ (Eq. 14) for the RS corpus.
  • LaMReDA (BiomedBERT large) with the relation representation $R_{F}$ (Eq. 14) for the AD corpus.
We stress that the selection of the models relies on the performance in the 5-fold cross-validation setup to avoid choosing based on the model performance on the original test set. Each model is trained 10 times with different seeds using the original splits (Tab. 1) of ReDReS and ReDAD correspondingly. The best model weights are saved based on the performance on the development set. Every trained model provides the predictions for the unannotated instances and the final silver labels are extracted through majority voting. The Distantly Supervised Relation Detection dataset for Rett Syndrome (DiSReDReS) contains 304,008 instances with 8,611 unique CUIs and 80 semantic types (Tab. 5). The Distantly Supervised Relation Detection dataset for Alzheimer’s Disease (DiSReDAD) comprises 13,608,175 instances with 53,750 unique CUIs and 82 semantic types (Tab. 5). As noisy labeling is inevitable in distantly supervised data and imposes challenges for knowledge extraction scenarios, the two extensive datasets can promote weakly supervised learning.
Data Sentences Instances CUIs1 S.T.1 Benchmark3
Binary Multi-Class
Micro Macro
DiSReDReS 28,622 304,008 8,611 80 91.53 75.1 75.19
DiSReDAD 1,301,429 13,608,175 53,750 82 88.99 80.56 80.69
  • 1
    The total number of unique CUIs.
  • 2
    Semantic Types.
  • 3
Table 11: DiSReDReS & DiSReDAD: Statistics and performance (%)
Dataset Labels - Type of relation
Positive Complex Negative No Relation
DiSReDAD 4,468,110 (32.8%) 5,755,884 (42.3%) 120,267 (0.9%) 3,263,914 (24%)
Table 12: DiSReDReS & DiSReDAD: Label Distribution
Weakly Supervised Setup. The task formulation remains the same as described in section 3 of the paper. The train sets of ReDReS and ReDAD are replaced by DiSReDReS and DiSReDAD, respectively. The development and test sets remain the same (Tab. 1). To provide a benchmark, we train the LaMReDA (BiomedBERT base) with the relation representation $R_{A}$ (Eq. 9) for 10 epochs utilizing the ADAM optimizer with learning rate 10-5. The batch is set to 32. The experiments are repeated 10 times with different seeds and the best scores are retained based on the performance on the development set.
In the supervised setup (Tab. 1), LaMReDA (BiomedBERT base) with the $R_{A}$ representation achieves 90.72% and 88.31% F1-score in the binary setup on ReDReS and ReDAD, respectively (Tab. 3). Multi-class macro F1-scores range from 74.52% (micro: 74.49%) to 77.34% (micro: 77.64%) for ReDReS and ReDAD, accordingly (Tab. 3). Table 5 presents the benchmark performance in the binary and multi-class setup for both datasets. Notably, the performance is improved in the weakly supervised setup, indicating the robustness of LaMReDA, when trained with noisy data, and highlighting the quality of the silver labels of DiSReDReS and DiSReDAD.
Appendix I Probing: Additional Experiments
Figure 12: ReDReS Probing (Multi-class setup, Macro evaluation) (BiomedBERT base): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
We use the same experimental setup as described in subsection 3.3 and the experiments are conducted in the 5-fold cross-validation setting. To provide an inclusive probing analysis on ReDReS, we incorporate additional probing results in this section. Fig. 8 and 9 present the experiments in the multi-class setup using BiomedBERT base. Additionally, aiming to explore the probing capabilities of BiomedBERT large, we include the results of further experiments in Fig. 10, 11, and 12. These experiments investigate the model’s performance in detecting semantic relations, comparing the representations and attention mechanisms at different layers and heads to understand how well the larger LM can discern complex relationships in the biomedical text.
Figure 13: ReDReS Probing (Multi-class setup, Micro evaluation) (BiomedBERT base): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
Figure 14: ReDReS Probing (Binary setup) (BiomedBERT large): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
Figure 15: ReDReS Probing (Multi-class setup, Macro evaluation) (BiomedBERT large): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
Figure 16: ReDReS Probing (Multi-class setup, Micro evaluation) (BiomedBERT large): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
Appendix J Detailed Annotation Guidelines
Task. Determine if there is any semantic relation between the two colored entities in the sentence.

### v3

@@ Appendix B Annotation Portal
Fig. 13 presents the annotation portal with an example from the ReDReS dataset. The annotator’s task is to identify the semantic relation between the two highlighted entities, classifying it as either a Positive Relation, Negative Relation, Complex Relation, or No Relation. If a sentence is considered uninformative or if there are errors in entity detection, type, or span, the annotator can remove the sentence or entities. Furthermore, the annotator is encouraged to provide feedback, including any additional text that can clarify or elaborate on the relationship between the entities. By providing this supplementary information, annotators can contribute to a richer and more nuanced understanding of the relations within the data.
Figure 13: Annotation portal: The annotator should define the semantic relation (Positive Relation, Negative Relation, Complex Relation, and No Relation) between the two highlighted entities: Rett syndrome and MECP2 gene. The annotator can remove the sentence if it is not informative and the entities if the entity detection or the type/span is incorrect. Additionally, the annotator can provide feedback, adding text that is useful for defining the relationship between the entities.
@@ Appendix C Dataset Instances
  • These techniques have thus provided morphological and functional brain alterations mapping of Alzheimer’s disease: on one hand grey matter atrophy first concerns the medial temporal lobe before extending to the temporal neocortex and then other neocortical areas; on the other hand, metabolic alterations are first located within the posterior cingulate cortex and then reach the temporoparietal area as well as the prefrontal cortex, especially in its medial part.
@@ Appendix D Localized Context Vector
Appendix E Baseline Performance: Lower Bound
To establish a baseline performance (lower bound) for comparison, in contrast to the human evaluation that serves as an upper bound, we randomly assign class labels to each instance of the test set based on the prior class distribution in the training set (Tab. II). This simulates a classifier with no ability to learn relations between entities. We repeat this experiment 1 million times for robustness and report the average F1-score.
Figure 14: ReDReS Probing (Multi-class setup, Macro evaluation) (BiomedBERT base): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Table IV). Second boundary: classifier with average attention scores across all layers as input.
Figure 15: ReDReS Probing (Multi-class setup, Micro evaluation) (BiomedBERT base): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Table IV). Second boundary: classifier with average attention scores across all layers as input.
Appendix F Probing: Additional Experiments
We use the same experimental setup as described in subsection 3.3 and the experiments are conducted in the 5-fold cross-validation setting. To provide an inclusive probing analysis on ReDReS, we incorporate additional probing results in this section. Fig. 14 and 15 present the experiments in the multi-class setup using BiomedBERT base. Additionally, aiming to explore the probing capabilities of BiomedBERT large, we include the results of further experiments in Fig. 16, 17, and 18. These experiments investigate the model’s performance in detecting semantic relations, comparing the representations and attention mechanisms at different layers and heads to understand how well the larger LM can discern complex relationships in biomedical text.
Figure 16: ReDReS Probing (Binary setup) (BiomedBERT large): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Table IV). Second boundary: classifier with average attention scores across all layers as input.
Figure 17: ReDReS Probing (Multi-class setup, Macro evaluation) (BiomedBERT large): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Table IV). Second boundary: classifier with average attention scores across all layers as input.
Figure 18: ReDReS Probing (Multi-class setup, Micro evaluation) (BiomedBERT large): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Table IV). Second boundary: classifier with average attention scores across all layers as input.
Appendix G Results with standard deviation
We present the results with the standard deviation when we utilize the official split of the data (Table II) and run the experiments 10 times with different seeds (Tables XI, XII, XIII and XIV).
TABLE XI: LaMEL Results (%) in binary setup utilizing BiomedBERT-base, BioLinkBERT-base, and BioGPT as backbone language model: Each cell shows the average F1-score from 10 runs (original test set) with the standard deviation.
Type1 ReDReS ReDAD
BiomedBERT2 BioLinkBERT2 BioGPT BiomedBERT2 BioLinkBERT2 BioGPT
A 90.25±0.7 90.68±0.9 89.43±1.1 86.73±1.7 87.71±0.8 86.14±1.6
D 90.47±0.8 90.35±0.9 88.52±0.9 86.29±1.0 87.35±1.3 85.28±1.9
E 90.61±0.6 90.25±0.6 89.48±0.8 86.03±1.0 87.05±1.2 87.42±0.8
H 89.68±1.3 90.73±0.6 89.21±0.9 87.29±1.0 87.56±1.1 85.55±2.0
@@ Appendix G Results with standard deviation
TABLE XII: LaMEL Results (%) in binary setup (BiomedBERT $\square$: base, $\blacksquare$:large): Each cell shows the average F1-score from 10 runs (original test set) with the standard deviation.
@@
F1$\square$ F1$\blacksquare$ F1$\square$ F1$\blacksquare$
A 90.25±0.7 90.88±0.7 86.73±1.7 88.90±0.7
B 90.29±0.5 90.73±0.7 86.89±1.1 88.22±1.1
C 90.51±0.5 90.71±0.6 87.49±0.9 88.57±1.1
D 90.47±0.8 91.03±0.6 86.29±1.0 88.22±0.9
E 90.61±0.6 90.54±0.7 86.03±1.0 88.74±1.4
F 90.48±1.1 90.88±0.4 87.27±0.9 89.44±0.5
G 90.32±0.9 90.35±0.7 86.97±1.1 89.12±0.5
H 89.68±1.3 90.13±1.4 87.29±1.0 88.77±1.2
CD2 86.20±1.1 89.14±1.0 88.92±0.5 88.56±0.8
  • 1
    Type of Relation Representation.
  • 2
TABLE XIII: LaMReDA and LaMReDM Results (%) in binary and multi-class setup (BiomedBERT $\square$: base,$\blacksquare$: large): Each cell shows the average F1-score from 10 runs (original test set) with the standard deviation.
Data Type1 Binary setup Multi-class setup
Micro Evaluation Macro Evaluation
LaMReDA LaMReDM LaMReDA LaMReDM LaMReDA LaMReDM
ReDReS A 90.72±0.4 90.74±0.7 90.42±1.0 90.71±0.5 74.49±1.3 73.96±0.8 74.36±1.0 74.35±0.8 74.52±1.3 73.66±1.1 74.30±0.9 72.81±1.3
B 90.40±0.4 90.28±0.9 90.47±0.7 90.06±0.6 74.27±1.1 73.72±1.3 74.26±1.3 73.57±0.9 74.32±1.1 73.65±1.1 74.38±1.2 73.11±1.2
C 90.85±0.7 90.75±0.8 90.51±0.5 89.14±0.8 74.93±1.2 73.54±1.7 74.31±1.1 73.69±1.5 74.96±1.1 73.44±1.9 74.10±0.9 73.49±0.7
D 90.55±0.4 90.93±0.7 90.61±0.4 90.53±0.4 73.61±0.6 73.50±1.1 73.02±0.9 73.21±0.8 73.71±0.8 73.70±1.2 73.24±0.5 73.90±1.0
E 89.57±0.8 89.43±0.6 89.57±0.8 89.43±0.6 73.73±0.9 73.68±1.0 73.73±0.9 73.68±1.0 73.95±0.7 74.01±0.8 73.95±0.7 74.01±0.8
F 90.48±0.4 90.62±0.8 90.41±0.3 90.43±0.8 72.86±1.1 73.82±0.9 73.51±0.9 73.33±0.9 72.62±0.7 73.94±1.2 73.32±1.1 74.50±0.8
G 90.78±0.7 90.76±0.9 90.49±0.5 89.47±0.9 74.33±1.1 73.63±1.3 74.05±1.2 73.22±1.0 74.57±0.7 73.31±1.7 74.21±1.2 73.87±1.0
H 90.91±0.5 90.45±0.7 90.29±0.8 89.99±0.6 74.43±0.8 73.62±1.3 73.59±0.5 73.36±0.8 74.48±0.5 73.65±1.4 73.90±0.7 73.42±1.6
I 90.86±0.4 90.47±0.6 90.62±0.6 89.55±0.8 74.75±0.9 73.30±0.7 74.29±0.7 74.00±1.5 74.80±0.8 73.26±1.1 73.88±0.8 73.91±1.1
J 89.43±1.0 89.65±0.7 89.53±0.9 89.89±0.5 73.75±0.8 74.28±1.0 73.47±0.7 74.43±1.0 74.05±1.1 75.06±0.9 73.52±0.8 74.70±1.3
K 90.10±0.6 90.05±0.5 89.70±0.4 89.64±0.7 74.43±0.8 74.07±1.2 74.47±0.8 74.38±0.6 74.44±0.9 74.23±1.5 74.30±0.7 74.02±0.9
L 89.60±0.8 89.85±0.8 89.82±0.7 90.33±0.6 73.35±1.8 74.32±0.7 73.68±1.0 73.90±0.9 73.16±1.8 74.04±0.6 73.66±1.1 73.52±1.2
M 90.81±0.8 90.07±0.6 90.01±0.7 89.75±0.7 74.21±0.9 74.29±0.7 73.96±0.9 74.32±1.0 74.03±0.8 74.36±0.7 73.72±0.7 73.95±1.1
N 90.73±0.7 90.60±0.8 90.72±0.8 90.63±0.8 74.55±1.2 73.83±1.3 73.86±0.9 73.38±1.6 74.66±1.3 73.97±1.0 74.13±1.0 73.53±0.9
O 90.90±0.5 90.50±1.0 90.90±0.5 90.50±1.0 73.99±1.1 73.77±1.2 73.99±1.1 73.77±1.2 73.83±1.2 73.62±1.1 73.83±1.2 73.62±1.1
P 89.72±0.7 90.31±0.4 89.13±0.7 90.30±0.6 73.37±1.0 73.87±1.1 73.57±1.0 74.41±1.0 73.48±1.1 74.66±1.3 73.54±1.0 73.52±1.1
CD2 87.42±0.5 88.93±0.9 87.76±0.7 88.10±0.7 73.09±1.0 75.04±1.6 74.15±0.8 75.35±1.2 73.64±0.8 74.94±1.6 74.38±1.5 75.44±1.8
ReDAD A 88.31±0.8 89.55±0.9 87.98±0.8 89.14±0.8 77.64±1.4 79.47±1.5 78.34±1.3 80.21±0.8 77.34±1.1 79.26±1.8 78.44±1.3 80.13±0.8
B 87.82±0.7 89.11±0.9 87.66±0.5 88.64±0.9 77.74±1.0 78.65±0.6 78.61±0.9 78.91±1.1 77.13±1.1 78.58±0.6 77.76±0.8 78.98±0.8
C 88.30±0.6 89.21±2.4 88.11±0.6 89.17±1.8 77.14±1.5 79.67±1.7 78.19±1.1 79.32±1.1 77.08±1.2 79.35±1.6 77.82±0.9 79.26±1.1
D 87.33±0.9 89.82±0.5 88.18±0.8 88.80±2.5 78.28±0.9 79.54±1.4 76.81±1.2 78.68±1.9 78.26±1.0 78.47±1.7 76.67±1.2 78.73±1.3
E 88.03±0.9 89.37±0.6 88.03±0.9 89.37±0.6 77.83±0.9 77.54±1.3 77.83±0.9 77.54±1.3 77.75±0.9 77.44±1.4 77.75±0.9 77.44±1.4
F 87.71±0.8 88.45±2.9 87.87±0.6 88.54±0.6 77.59±1.7 79.74±1.4 76.94±2.2 79.68±1.1 77.35±1.3 79.19±1.6 76.95±1.7 79.21±1.1
G 88.17±0.8 89.83±0.8 88.22±0.8 89.55±0.9 77.83±1.5 79.39±1.0 78.13±2.2 79.09±1.8 77.5±1.3 79.12±1.0 77.88±2.4 78.74±1.5
H 88.01±0.6 88.76±1.5 87.73±0.6 88.99±0.8 77.12±1.5 79.57±0.9 78.11±0.7 79.40±0.8 77.09±1.6 78.65±1.3 78.14±0.5 79.18±0.6
I 87.56±2.4 88.05±3.9 87.86±1.6 89.45±0.8 77.77±1.4 79.23±1.1 78.40±1.1 78.99±0.9 77.11±1.2 79.05±1.1 78.49±1.1 78.78±1.1
J 87.99±0.4 88.89±1.5 87.79±0.4 89.06±0.9 77.69±0.8 78.40±1.1 77.14±3.0 78.40±0.9 77.71±0.7 77.94±1.2 76.59±2.7 78.18±0.8
K 88.36±0.7 89.33±0.6 88.01±0.6 89.05±0.8 78.30±1.1 78.25±0.8 78.54±0.9 77.73±1.9 78.11±1.4 78.49±0.9 78.29±0.7 77.42±1.3
L 88.25±0.6 89.25±0.6 87.87±0.6 89.13±0.8 78.48±0.9 78.94±0.6 77.88±0.6 77.91±1.0 78.52±0.9 78.16±0.9 77.85±0.6 77.37±1.4
M 88.42±0.5 89.57±0.5 88.12±0.7 89.40±0.6 77.00±1.8 78.85±0.8 78.02±1.2 77.12±0.7 76.62±2.0 79.66±1.0 78.02±1.2 77.08±1.0
N 87.98±0.8 88.94±0.9 88.08±0.5 89.47±0.5 78.21±0.8 78.92±0.9 78.03±1.1 78.78±1.6 78.07±0.9 77.91±0.7 77.74±1.0 78.60±1.4
O 88.27±0.8 87.71±2.8 88.27±0.8 87.71±2.8 77.08±1.4 78.96±1.6 77.08±1.4 78.96±1.8 76.78±1.7 79.03±1.6 76.78±1.7 79.03±1.6
P 88.02±1.0 88.86±1.2 88.33±1.0 89.51±1.0 78.43±0.8 79.38±1.4 78.44±0.7 79.12±1.1 78.37±0.9 79.44±1.2 78.04±1.2 78.91±1.1
CD2 88.40±0.5 89.33±0.8 89.01±0.9 89.16±0.6 73.69±1.8 74.29±1.8 72.67±2.1 72.81±1.4 74.13±1.9 74.82±1.7 72.91±1.7 73.76±1.6
  • 1
    Type of Relation Representation.
  • 2
TABLE XIV: LaMReDA Results (%) in binary and multi-class setup utilizing BiomedBERT-base, BioLinkBERT-base, and BioGPT as backbone language model: Each cell shows the average F1-score from 10 runs (original test set) with the standard deviation.
Data Type1 Binary setup Multi-class setup
Micro Evaluation Macro Evaluation
ReDReS A 90.72±0.4 91.15±0.4 90.15±0.5 74.49±1.3 74.06±0.9 72.22±1.2 74.52±1.3 74.11±1.0 71.80±1.2
D 90.55±0.4 90.35±1.1 89.76±0.9 73.61±0.6 74.12±0.9 71.42±0.7 73.71±0.8 74.12±0.7 72.48±0.9
E 89.57±0.8 89.98±0.6 89.49±0.6 73.73±0.9 73.22±0.6 70.17±1.1 73.95±0.7 73.38±1.1 69.97±0.6
F 90.48±0.4 90.72±0.5 89.68±0.8 72.86±1.1 74.22±0.8 71.36±1.1 72.62±0.7 74.13±0.7 72.00±1.4
G 90.78±0.7 90.57±0.6 89.93±0.7 74.33±1.1 74.00±1.5 72.68±1.1 74.57±0.7 74.32±1.6 72.48±0.7
O 90.90±0.5 90.75±0.9 89.51±0.9 73.99±1.1 74.00±1.5 66.48±4.9 73.83±1.2 73.69±1.7 67.43±2.4
ReDAD A 88.31±0.8 88.73±1.0 87.49±0.9 77.64±1.4 78.40±0.8 76.04±1.4 77.34±1.1 77.68±0.8 76.45±0.8
D 87.33±0.9 88.74±0.9 87.56±1.0 78.28±0.9 78.27±0.7 76.49±0.6 78.26±1.0 77.89±1.4 76.25±1.1
E 88.03±0.9 88.43±0.7 87.71±0.7 77.83±0.9 78.96±0.9 75.35±1.1 77.75±0.9 78.00±0.7 75.36±1.1
F 87.71±0.8 88.62±0.9 87.63±0.5 77.59±1.7 77.43±0.8 76.34±1.3 77.35±1.3 77.48±0.8 75.47±1.1
G 88.17±0.8 88.72±0.8 87.55±0.8 77.83±1.5 77.76±1.0 76.35±1.1 77.50±1.3 77.33±1.1 76.61±0.6
O 88.27±0.8 88.58±0.6 86.19±1.5 77.08±1.4 76.89±0.9 71.69±3.2 76.78±1.7 76.62±1.3 70.57±3.4
@@
Appendix H Detailed Annotation Guidelines